OpenStudy (hba):

If a complex number z=x+iy then in terms of conjugate coordinates x can be written as :

OpenStudy (hba):

@hartnn

hartnn (hartnn):

the question is 'x can be.....' or 'z can be....' ?

OpenStudy (hba):

x can be

hartnn (hartnn):

then i need options...

OpenStudy (hba):

z-zbar,z+zbar/2,none of the above

hartnn (hartnn):

ohh....so find z bar first, what u get ?

OpenStudy (hba):

How do i do that is it underroot(x^2+y^2)

hartnn (hartnn):

ok, if z = a+bi x bar = conjugate = a-bi got this ?

hartnn (hartnn):

*z bar

OpenStudy (hba):

Ok

hartnn (hartnn):

now what is z bar here ?

OpenStudy (hba):

zbar=(x-yi)

hartnn (hartnn):

now to eliminate y here, we add z and z bar what u get ?

OpenStudy (hba):

z+zbar=(x+yi)+(x-yi)=2x

hartnn (hartnn):

so whats x ?

OpenStudy (hba):

z+zbar/2 .Right Thanks :)

hartnn (hartnn):

welcome ^_^